# Around the Merino–Welsh conjecture: improving Jackson’s inequality

### v1

@@ Abstract.
The Merino-Welsh conjecture states that for a graph $G$ without loops and bridges we have
$\max(T_{G}(2,0),T_{G}(0,2))\geq T_{G}(1,1).$
Later Jackson proved that for any matroid $M$ without loop and coloop we have
$T_{M}(3,0)T_{M}(0,3)\geq T_{M}(1,1)^{2}.$
The value $3$ in this statement was improved to $2.9242$ by Beke, Csáji, Csikvári and Pituk. In this paper, we further improve on this result by showing that
$T_{M}(2.355,0)T_{M}(0,2.355)\geq T_{M}(1,1)^{2}.$
We also prove that the Merino–Welsh conjecture is true for matroids $M$, where all circuits of $M$ and its dual $M^{*}$ have length between $\ell$ and $(\ell-2)^{4}$ for some $\ell\geq 6$.
The research was supported by the MTA-Rényi Counting in Sparse Graphs ”Momentum” Research Group, and by Dynasnet European Research Council Synergy project – grant number ERC-2018-SYG 810115.
@@ 1. Introduction
$\max\left(\alpha(G),\alpha^{*}(G)\right)\geq\tau(G),$
@@
Conde and Merino [5] proposed "additive" and "multiplicative" versions of this conjecture:
$T_{G}(2,0)+T_{G}(0,2)$ $\geq 2T_{G}(1,1)$
$T_{G}(2,0)T_{G}(0,2)$ $\geq T_{G}(1,1)^{2},$
@@
Furthermore, let $x_{0}$ be the largest root of the polynomial $x^{3}-9(x-1)$. ($x_{0}\approx 2.22668...$) Then for $0\leqslant a<x_{0}$ there are infinitely many matroids $M$ without loops and coloops for which
@@
For any matroid M without loops and coloops and $a\geq 2.355$,
$T_{M}(a,0)T_{M}(0,a)\geq T_{M}(1,1)^{2}.$
We may also study which matroid classes satisfy the product version of the Merino–Welsh conjecture. The following theorem is motivated by the fact that paving matroids satisfy the Merino–Welsh conjecture. A matroid of rank $r$ is a paving matroid if all circuits have length $r$ or $r+1$.
@@
Suppose that there exists an $\ell\geq 6$ such that all circuits of the matroid $M$ and its dual $M^{*}$ have length between $\ell$ and $(\ell-2)^{4}$. Then
$T_{M}(2,0)T_{M}(0,2)\geq T_{M}(1,1)^{2}.$
The proof of Theorem 1.3 and Theorem 1.4 are based on the theory of the permutation Tutte polynomial developed in the paper [1]. While Theorem 1.4 does not imply that paving matroids satisfy the Merino–Welsh conjecture, one can prove this fact by modifying the proof of Theorem 1.4.
Notations. Throughout the paper $G=(V,E)$ is an arbitrary graph and $H=(A,B,E)$ is a bipartite graph. $K_{a,b}$ denotes the complete bipartite graph with parts of size $a$ and $b$. $S_{k}$ denotes the star graph on $k$ vertices. For a vertex $v$ the degree of $v$ is denoted by $d_{v}$. $N_{H}(v)$ denotes the set of neighbors of $v$. If $H$ is clear from the context, then we simply write $N(v)$.
@@
  • In the next section we introduce the basic concepts from matroid theory that we will use, and revisit the theory of permutation Tutte polynomial $\widetilde{T}_{H}(x,y)$ developed in the paper [1].
  • In Section 3 we compute the growth constant $\lim_{n\to\infty}\widetilde{T}_{H_{n}}(x,0)^{1/n}$ for some family of bipartite graphs. While this section is not necessary for the proofs of Theorem 1.3 and 1.4 it provides an important intuition to study the permutation Tutte polynomial.
@@
  • In Section 7 we end the paper some conjectures.
@@ 2. Preliminaries
This section collects the necessary tools from matroid theory together, and recall some of the basic facts from the theory of permutation Tutte polynomial.
@@ 2.1. Tutte polynomial and matroids
with $k(A)$ denoting the number of connected components of the graph $(V,A)$, see [22]. There is a vast amount of literature on the properties of the Tutte polynomial and its applications, see for instance, [3, 6, 8, 23], or the book [9].
The Tutte polynomial naturally extends to matroids. Recall that a matroid $M$ is a pair $(E,\mathcal{I})$ such that $\mathcal{I}\subseteq 2^{E}$, called independent sets, satisfies the axioms (i) $\emptyset\in\mathcal{I}$, (ii) if $A^{\prime}\subseteq A\in\mathcal{I}$, then $A^{\prime}\in\mathcal{I}$, and (iii) if $A,B\in\mathcal{I}$ such that $|B|<|A|$, then there exists an $x\in A\setminus B$ such that $B\cup\{x\}\in\mathcal{I}$. Given a set $S\subseteq E$, the maximal independent subsets of $S$ all have the same cardinality, and this cardinality is called the rank of the matroid, denoted by $r(S)$. The maximum size independent sets of $M$ are called bases, and their set is denoted by $\mathcal{B}(M)$. The dual of a matroid $M$ is the matroid $M^{*}$ whose bases are $\{E\setminus B\ |\ B\in\mathcal{B}(M)\}$. For further details on matroids, see for instance [20]
Given a graph $G=(V,E)$, the edge sets of the spanning forests of $G$ form the independent sets of a matroid $M_{G}$ called the cycle matroid of $G$. If $G$ is connected, then the basis of $M_{G}$ are the spanning trees of $G$. One can define the Tutte polynomial of a matroid as
@@
where $r(S)$ is the rank of a set $S\subseteq E$. When $M=M_{G}$, then $T_{M_{G}}(x,y)=T_{G}(x,y)$. A loop in a matroid $M$ is an element $x\in E$ such that $r(\{x\})=0$, that is, $\{x\}\notin\mathcal{I}$, and a coloop is an element that is a loop in the dual $M^{*}$ of the matroid $M$. Equivalently, a coloop is an element that is in every base of $M$. For a cycle matroid $M_{G}$, loops correspond to loop edges and coloops correspond to bridges in the graph $G$.
Hence it was suggested that the inequalities
@@
may hold true for all matroids $M$ without loops and coloops. (These versions appear explicitly in [10], but were treated much earlier without explicitly calling them conjectures.) Note that for general matroids, all these versions are equivalent in the following sense: if one of them is true for all matroids, then the others are also true for all matroids. Applying the maximum version to $M\oplus M^{*}$ with $M^{*}$ being the dual of $M$ leads to the multiplicative version of the conjecture. (Here $M\oplus N$ denotes the direct sum of the matroids $M$ and $N$.)
@@ 2.2. Permutation Tutte polynomial
The proof of Theorem 1.3 heavily relies on the theory of permutation Tutte polynomial. The idea is that Tutte polynomial $T_{G}(x,y)$ can be written as a sum of permutation Tutte polynomials $\widetilde{T}_{H_{j}}(x,y)$ for certain bipartite graphs $H_{j}$. As a consequence certain inequalities valid for the permutation Tutte polynomial transfers to the Tutte polynomial.
@@
where the maximum over an empty set is set to be $\infty$- ∞. Similarly, we say that vertex $j\in B$ is externally active if
@@
$\widetilde{T}_{H}(x,y)=\frac{1}{m!}\sum_{\pi\in S_{m}}x^{\mathrm{ia}(\pi)}y^{\mathrm{ea}(\pi)}.$
We will call $\widetilde{T}_{H}(x,y)$ the permutation Tutte polynomial of $H$.
@@
Let $G$ be a connected graph with $m$ edges. Label the edges with $1,2,\dots,m$ arbitrarily. In the case of a spanning tree $T$ of $G$, let us call an edge $e\in E(T)$ internally active if $e$ has the largest label among the edges in the cut determined by $T$ and $e$ by removing $e$ from $T$. Let us call an edge $e\notin E(T)$ externally active if $e$ has the largest label among the edges in the cycle determined by $T$ and $e$ by adding $e$ to $T$. Let $\mathrm{ia}(T)$ and $\mathrm{ea}(T)$ be the number of internally and externally active edges, respectively. Then
@@
where the summation goes for all spanning trees of $G$.
Theorem 2.2 was originally a definition for the Tutte polynomial [22]. This characterization of the Tutte polynomial immediately shows that the coefficients of the Tutte polynomial are non-negative. In this theorem, we are restricted to the same labelling of the edges for all spanning trees. For those who have never seen this definition before, it might be very surprising that the Tutte polynomial is independent of the actual choice of the labelling.
@@
The local basis exchange graph $H[T]$ of a graph $G=(V,E)$ with respect to a spanning tree $T$ is defined as follows. The graph $H[T]$ is a bipartite graph whose vertices are the edges of $G$. One bipartite class consists of the edges of $T$, the other consists of the edges of $E\setminus T$, and we connect a spanning tree edge $e$ with a non-edge $f$ if $f$ is in the cut determined by $e$ and $T$, equivalently, $e$ is in the cycle determined by $f$ and $T$.
Clearly, this definition works for general matroids and their basis: if $A$ is a basis, then $e\in A$ and $f\in E\setminus A$ are adjacent in the bipartite graph $H[A]$ if $A-e+f$ is again a basis.
Figure 1 depicts a graph $G$ with a spanning tree $T$ and the bipartite graph $H[T]$ obtained from $T$.
For a fixed labelling of the edges of $G$, we get a labelling of the vertices of $H[T]$, and the internally (externally) active edges of $G$ correspond to internally (externally) active vertices of $H[T]$, so the two definitions of internal and external activity are compatible. The following lemma is crucial for us, so we even included its proof.
11112222333344446666555577778888
11112222333355557777444466668888
@@
Let $G$ be a graph. For each spanning tree $T$ of $G$, let $H[T]$ be the local basis exchange graph with respect to $T$. Then
$T_{G}(x,y)=\sum_{T\in\mathcal{T}(G)}\widetilde{T}_{H[T]}(x,y),$
where the sum is over the set of spanning trees $\mathcal{T}(G)$ of $G$.
@@
For a fixed spanning tree $T$ and a permutation $\pi$ of the edges, the internally and externally active edges correspond to the internally and externally active vertices of $H[T]$. Hence
$T_{G}(x,y)=\sum_{T\in\mathcal{T}(G)}x^{\mathrm{ia}_{H[T]}(\pi)}y^{\mathrm{ea}_{H[T]}(\pi)}.$
@@
$T_{G}(x,y)$ $=\frac{1}{m!}\sum_{\pi\in S_{m}}T_{G}(x,y)$
$=\frac{1}{m!}\sum_{\pi\in S_{m}}\sum_{T\in\mathcal{T}(G)}x^{\mathrm{ia}_{H[T]}(\pi)}y^{\mathrm{ea}_{H[T]}(\pi)}$
$=\sum_{T\in\mathcal{T}(G)}\frac{1}{m!}\sum_{\pi\in S_{m}}x^{\mathrm{ia}_{H[T]}(\pi)}y^{\mathrm{ea}_{H[T]}(\pi)}$
$=\sum_{T\in\mathcal{T}(G)}\widetilde{T}_{H[T]}(x,y).$
@@
The local basis exchange graph $H[T]$ has an isolated vertex if and only if $G$ contains a bridge or a loop. Furthermore, $H[T]$ is connected if and only $G$ is $2$-connected.
The following lemma enables us to study quadratic inequalities of the Tutte polynomial.
@@
Let $x_{0},x_{1},x_{2},y_{0},y_{1},y_{2}\geq 0$. Suppose that for any bipartite graph $H$, we have
$\widetilde{T}_{H}(x_{1},y_{1})\widetilde{T}_{H}(x_{2},y_{2})\geq\widetilde{T}_{H}(x_{0},y_{0})^{2}.$
Then for any graph $G$, we have
$T_{G}(x_{1},y_{1})T_{G}(x_{2},y_{2})\geq T_{G}(x_{0},y_{0})^{2}.$
@@
For us the following version of the transfer lemma will be a bit more convenient. If for any basis $A$ of a matroid $M$, the local basis exchange graph $H[A]$ satisfies that
$\widetilde{T}_{H[A]}(x_{1},y_{1})\widetilde{T}_{H[A]}(x_{2},y_{2})\geq\widetilde{T}_{H[A]}(x_{0},y_{0})^{2},$
then
$T_{M}(x_{1},y_{1})T_{M}(x_{2},y_{2})\geq T_{M}(x_{0},y_{0})^{2}.$
The proof of this claim is exactly the same as of the transfer lemma.
@@
Let $H$ be an arbitrary bipartite graph, and let $d_{i}$ be the degree of vertex $i$. Suppose that $0\leq x\leq 1$ and $y\geq 1$ or $0\leq y\leq 1$ and $x\geq 1$. Then
$\widetilde{T}_{H}(x,y)\geq\prod_{i\in A}\left(1+\frac{x-1}{d_{i}+1}\right)\cdot\prod_{j\in B}\left(1+\frac{y-1}{d_{j}+1}\right).$
The proof of Lemma 2.8 relies on following inequality of Harris that is also a special case of the FKG-inequality [11].
@@
Suppose that $\mu=\mu_{1}\otimes\dots\otimes\mu_{N}$ is a product measure on a set $S_{1}\times S_{2}\times\dots\times S_{N}$, and $X_{1},\dots,X_{t}$ are non-negative monotone increasing functions in the sense that if $x_{i}\geq x_{i}^{\prime}$ for $i=1,\dots,N$, then for $1\leq j\leq t$ we have
$X_{j}(x_{1},\dots,x_{N})\geq X_{j}(x_{1}^{\prime},\dots,x_{N}^{\prime}).$
@@
$\mathbb{E}_{\mu}\left[\prod_{j=1}^{t}X_{j}\right]\geq\prod_{j=1}^{t}\mathbb{E}_{\mu}[X_{j}].$
Furthermore, if $X$ is monotone increasing and $Y$ is monotone decreasing, then
$\mathbb{E}[XY]\leq\mathbb{E}[X]\mathbb{E}[Y].$
In this paper, we will use the Harris inequality for both $[0,1]^{N}$ and the set $\{0,1\}^{A}$ for some set $A$.
In what follows, we repeatedly use the following crucial idea to express $\widetilde{T}_{H}(x,y)$. We can create a random ordering of the vertices of $H$ as follows: for each vertex $i$ we choose a uniform random number $x_{i}$ from the interval $[0,1]$. The numbers $x_{i}$ then determine an ordering of the edges. The probability that two numbers are equal is 00.
@@
In what follows we do a little trick. For $i\in A$ we generate $x_{i}\sim U(0,1)$ as before, but for $j\in B$ we actually first generate a uniformly random number $y_{j}$ from $[0,1]$ and let $x_{j}=1-y_{j}$. The role of this trick will be apparent soon.
@@
$X_{i}(x_{i},\{y_{j}\}_{j\in B})=\left\{\begin{array}[]{ll}x&\mbox{if}\ \max_{j\in N_{H}(i)}(1-y_{j})\leq x_{i},\\ 1&\mbox{if}\ \max_{j\in N_{H}(i)}(1-y_{j})>x_{i}.\end{array}\right.$

### v2

@@ Abstract.
The Merino-Welsh conjecture states that for a graph $G$ without loops and bridges the Tutte polynomial $T_{G}(x,y)$ satisfies the inequality
$\max(T_{G}(2,0),T_{G}(0,2))\geqslant T_{G}(1,1).$
Later Jackson proved that for any matroid $M$ without loops and coloops we have
$T_{M}(3,0)T_{M}(0,3)\geqslant T_{M}(1,1)^{2}.$
The value $3$ in this statement was improved to $2.9243$ by Beke, Csáji, Csikvári and Pituk. In this paper, we further improve on this result by showing that
$T_{M}(2.355,0)T_{M}(0,2.355)\geqslant T_{M}(1,1)^{2}.$
We also prove that the Merino–Welsh conjecture is true for matroids $M$, where all circuits of $M$ and its dual $M^{*}$ have length between $\ell$ and $(\ell-2)^{2}(\ell^{2}-4\ell+2)$ for some $\ell\geqslant 4$.
The research was supported by the MTA-Rényi Counting in Sparse Graphs ”Momentum” Research Group, by the Hungarian National Research, Development and Innovation Office, Advanced grant 153378, and by Dynasnet European Research Council Synergy project – grant number ERC-2018-SYG 810115.
@@ 1. Introduction
$\max\left(\alpha(G),\alpha^{*}(G)\right)\geqslant\tau(G),$
@@
Conde and Merino [5] proposed “additive” and “multiplicative” versions of this conjecture:
$T_{G}(2,0)+T_{G}(0,2)$ $\geqslant 2T_{G}(1,1),$
$T_{G}(2,0)T_{G}(0,2)$ $\geqslant T_{G}(1,1)^{2},$
@@
Furthermore, let $x_{0}$ be the largest root of the polynomial $x^{3}-9(x-1)$. ($x_{0}\approx 2.22668$...) Then for any $0\leqslant a<x_{0}$ there are infinitely many matroids $M$ without loops and coloops for which
@@
For any matroid M without loops and coloops and $a\geqslant 2.355$,
$T_{M}(a,0)T_{M}(0,a)\geqslant T_{M}(1,1)^{2}.$
We may also study which matroid classes satisfy the product version of the Merino–Welsh conjecture. The following theorem is motivated by the fact that paving matroids satisfy the Merino–Welsh conjecture [4, 10]. A matroid of rank $r$ is a paving matroid if all circuits have length $r$ or $r+1$.
@@
Suppose that there exists an $\ell\geqslant 4$ such that all circuits of the matroid $M$ and its dual $M^{*}$ have length between $\ell$ and $(\ell-2)^{2}(\ell^{2}-4\ell+2)$. Then
$T_{M}(2,0)T_{M}(0,2)\geqslant T_{M}(1,1)^{2}.$
The proofs of Theorem 1.3 and Theorem 1.4 are based on the theory of the permutation Tutte polynomials developed in the paper [1]. While Theorem 1.4 does not imply that paving matroids satisfy the Merino–Welsh conjecture, one can prove this fact by modifying the proof of Theorem 1.4.
Notation. Throughout the paper $G=(V,E)$ is an arbitrary graph and $H=(A,B,E)$ is a bipartite graph. $K_{a,b}$ denotes the complete bipartite graph with parts of size $a$ and $b$. $S_{k}$ denotes the star graph on $k$ vertices, that is, $S_{k}=K_{1,k-1}$. For a vertex $v$ the degree of $v$ is denoted by $d_{v}$. $N_{H}(v)$ denotes the set of neighbors of $v$. If $H$ is clear from the context, then we simply write $N(v)$.
@@
  • In the next section, we introduce the basic concepts from matroid theory that we will use, and revisit the theory of the permutation Tutte polynomial $\widetilde{T}_{H}(x,y)$ developed in the paper [1].
  • In Section 3 we compute the growth constant $\lim_{n\to\infty}\widetilde{T}_{H_{n}}(x,0)^{1/n}$ for some families of bipartite graphs. While this section is not necessary for the proofs of Theorem 1.3 and 1.4 it provides an important intuition concerning the permutation Tutte polynomial.
@@
  • In Section 7 we end the paper with some conjectures.
@@ 2. Preliminaries
This section collects the necessary tools from matroid theory, and recalls some of the basic facts from the theory of the permutation Tutte polynomial.
@@ 2.1. Tutte polynomial and matroids
with $k(A)$ denoting the number of connected components of the graph $(V,A)$ and $v(G)$ denotes the number of vertices of $G$, see [22]. There is a vast literature on the properties of the Tutte polynomial and its applications, see for instance, [3, 6, 8, 23], or the book [9].
The Tutte polynomial naturally extends to matroids. Recall that a matroid $M$ is a pair $(E,\mathcal{I})$ such that $\mathcal{I}\subseteq 2^{E}$, the collection of independent sets, satisfies the axioms
  • (i)
    $\emptyset\in\mathcal{I}$,
  • (ii)
    if $A^{\prime}\subseteq A\in\mathcal{I}$, then $A^{\prime}\in\mathcal{I}$,
  • (iii)
    if $A,B\in\mathcal{I}$ such that $|B|<|A|$, then there exists an $x\in A\setminus B$ such that $B\cup\{x\}\in\mathcal{I}$.
It turns out that given a set $S\subseteq E$, the maximal independent subsets of $S$ all have the same cardinality. This cardinality is called the rank of the set $S$, and is denoted by $r(S)$. The maximum size independent sets of $M$ are called bases, and the set of bases is denoted by $\mathcal{B}(M)$. The dual of a matroid $M$ is the matroid $M^{*}$ whose bases are $\{E\setminus B\ |\ B\in\mathcal{B}(M)\}$. A circuit of a matroid is a minimal dependent set $A$, that is, each proper subset of $A$ is an independent set, but $A$ itself is not in $\mathcal{I}$. For further details on matroids, see for instance [20].
This paper assumes very little familiarity with matroid theory, and we offer the following two families of examples for those unfamiliar with the basics.
Let $U_{r,n}$ denote the uniform matroid of rank $r$ on $n$ elements: here the set $\mathcal{I}$ consists of the subsets of $\{1,2,\dots,n\}$ that have size at most $r$. Clearly, $\mathcal{I}$ satisfies all three axioms. Here the bases are the subsets of $\{1,2,\dots,n\}$ that have size exactly $r$. The circuits are the sets of size $r+1$, and the dual matroid is $U_{n-r,n}$.
Another important family of matroids are the so-called cycle matroids. Given a graph $G=(V,E)$ we can consider the matroid $M_{G}$ on the ground set $E$, where $A\subseteq E$ is independent if it does not contain a cycle, that is, it is a forest. It is clear that it satisfies (i) and (ii), and a little inspection shows that it also satisfies (iii). This is called the cycle matroid of $G$, and is denoted by $M_{G}$. One can see that if $G$ is connected, then the bases of $M_{G}$ are the spanning trees of $G$. In general, the rank of a set $A\subseteq E$ is v(G)k(A)v(G)-k(A), where $k(A)$ is the number of connected components of the subgraph $(V,A)$. A circuit of $M_{G}$ corresponds to the edge set of a cycle of $G$. We believe that understanding uniform matroids and cycle matroids is sufficient to have a good grasp on the content of this paper.
One can define the Tutte polynomial of a matroid as
@@
where $r(S)$ is the rank of a set $S\subseteq E$. When $M=M_{G}$, then $T_{M_{G}}(x,y)=T_{G}(x,y)$. A loop in a matroid $M$ is an element $x\in E$ such that $r(\{x\})=0$, that is, $\{x\}\notin\mathcal{I}$, and a coloop is an element that is a loop in the dual $M^{*}$ of the matroid $M$. Equivalently, a coloop is an element that is in every base of $M$. For a cycle matroid $M_{G}$, loops correspond to loop edges, and coloops correspond to bridges in the graph $G$.
Hence, it was suggested that the inequalities
@@
may hold true for all matroids $M$ without loops and coloops. (These versions appear explicitly in [10], but were treated much earlier without explicitly calling them conjectures.) Note that for general matroids, all these variants are equivalent in the following sense: if one of them is true for all matroids, then the others are also true for all matroids. Applying the maximum version to $M\oplus M^{*}$ with $M^{*}$ being the dual of $M$ leads to the multiplicative version of the conjecture. (Here $M\oplus N$ denotes the direct sum of the matroids $M$ and $N$.)
@@ 2.2. Permutation Tutte polynomial
The proof of Theorem 1.3 heavily relies on the theory of the permutation Tutte polynomial. The idea is that the Tutte polynomial $T_{G}(x,y)$ can be written as a sum of the permutation Tutte polynomials $\widetilde{T}_{H_{j}}(x,y)$ for certain bipartite graphs $H_{j}$. As a consequence certain inequalities valid for the permutation Tutte polynomial transfer to the Tutte polynomial.
@@
where the maximum over an empty set is set to be -$\infty$. Similarly, we say that vertex $j\in B$ is externally active if
@@
$\widetilde{T}_{H}(x,y)=\frac{1}{m!}\sum_{\pi\in S_{m}}x^{\mathrm{ia}(\pi)}y^{\mathrm{ea}(\pi)},$
where $S_{m}$ denotes the set of all permutations on $m$ elements. We will call $\widetilde{T}_{H}(x,y)$ the permutation Tutte polynomial of $H$.
@@
Let $G$ be a connected graph with $m$ edges. Label the edges with $1,2,\dots,m$ arbitrarily. In the case of a spanning tree $T$ of $G$, let us call an edge $e\in E(T)$ internally active if $e$ has the largest label among the edges $f^{\prime}\in E(G)$ that are going between the connected components of the graph TeT-e. Let us call an edge $f\notin E(T)$ externally active if $e$ has the largest label among the edges in the cycle determined by $T$ and $f$ by adding $f$ to $T$. Let $\mathrm{ia}(T)$ and $\mathrm{ea}(T)$ be the number of internally and externally active edges, respectively. Then
@@
where the summation runs over all spanning trees of $G$.
Theorem 2.2 was originally a definition for the Tutte polynomial [22]. This theorem or definition naturally extends to matroids even though Tutte’s original paper only concerned graphs. This characterization of the Tutte polynomial immediately shows that the coefficients of the Tutte polynomial are non-negative. In this theorem, we are restricted to the same labeling of the edges for all spanning trees. For those who have never seen this definition before, it might be very surprising that the Tutte polynomial is independent of the actual choice of the labeling.
@@
The local basis exchange graph $H[T]$ of a graph $G=(V,E)$ with respect to a spanning tree $T$ is defined as follows. The graph $H[T]$ is a bipartite graph whose vertices are the edges of $G$. One bipartite class consists of the edges of $T$, the other consists of the edges of $E\setminus T$, and we connect a spanning tree edge $e$ with a non-edge $f$ if $f$ is in the cut determined by $e$ and $T$, equivalently, $e$ is in the cycle determined by $f$ and $T$. This is also equivalent with Te+fT-e+f being a spanning tree of $G$ again.
Clearly, this definition works for general matroids $M=(E,\mathcal{I})$ and their basis. If $A$ is a basis, then let $H[A]$ be the bipartite graph with $V(H[A])=E$, where one part consists of the elements of $A$, the other part consists of $E\setminus A$, and $e\in A$ and $f\in E\setminus A$ are adjacent in the bipartite graph $H[A]$ if Ae+fA-e+f is again a basis.
Figure 1 shows a graph $G$ with a spanning tree $T$ and the bipartite graph $H[T]$ obtained from $T$.
For a fixed labeling of the edges of $G$, we get a labeling of the vertices of $H[T]$, and the internally (externally) active edges of $G$ correspond to internally (externally) active vertices of $H[T]$, so the two definitions of internal and external activity are compatible. The following lemma is crucial for us, so we even included its proof.
1122334466557788
1122335577446688
@@
Let $M$ be a matroid. For each basis $A$ of $M$, let $H[A]$ be the local basis exchange graph with respect to $A$. Then
$T_{M}(x,y)=\sum_{A\in\mathcal{B}(M)}\widetilde{T}_{H[A]}(x,y),$
where the sum runs over the set of bases $\mathcal{B}(M)$ of $M$.
@@
For a fixed basis $A$ and a permutation $\pi$ of the edges, the internally and externally active edges correspond to the internally and externally active vertices of $H[A]$. Hence
$T_{M}(x,y)=\sum_{A\in\mathcal{B}(M)}x^{\mathrm{ia}_{H[A]}(\pi)}y^{\mathrm{ea}_{H[A]}(\pi)}.$
@@
$T_{M}(x,y)$ $=\frac{1}{m!}\sum_{\pi\in S_{m}}T_{M}(x,y)$
$=\frac{1}{m!}\sum_{\pi\in S_{m}}\sum_{A\in\mathcal{B}(M)}x^{\mathrm{ia}_{H[A]}(\pi)}y^{\mathrm{ea}_{H[A]}(\pi)}$
$=\sum_{A\in\mathcal{B}(M)}\frac{1}{m!}\sum_{\pi\in S_{m}}x^{\mathrm{ia}_{H[A]}(\pi)}y^{\mathrm{ea}_{H[A]}(\pi)}$
$=\sum_{A\in\mathcal{B}(M)}\widetilde{T}_{H[A]}(x,y).$
@@
The local basis exchange graph $H[A]$ has an isolated vertex if and only if $M$ contains a loop or a coloop. Furthermore, in case of a graph $G$ and a spanning tree $T$, the graph $H[T]$ is connected if and only if $G$ is $2$-connected.
The following lemma is a variant of the transfer lemma [1] with the exact same proof. It enables us to study certain quadratic inequalities of the Tutte polynomial.
@@
Let $M$ be a matroid and $x_{0},x_{1},x_{2},y_{0},y_{1},y_{2}\geqslant 0$. Suppose that for any basis $A$ of the matroid $M$, the local basis exchange graph $H[A]$ satisfies
$\widetilde{T}_{H[A]}(x_{1},y_{1})\widetilde{T}_{H[A]}(x_{2},y_{2})\geqslant\widetilde{T}_{H[A]}(x_{0},y_{0})^{2},$
then
$T_{M}(x_{1},y_{1})T_{M}(x_{2},y_{2})\geqslant T_{M}(x_{0},y_{0})^{2}.$
@@
In this paper, we aim to prove the inequality
$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geqslant\widetilde{T}_{H}(1,1)^{2}$
for various values of $x$. Observe that $\widetilde{T}_{H}(1,1)=1$, a direct consequence of the definition. So what we really need to prove is
$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geqslant 1.$
@@
Let $H=(A,B,E)$ be an arbitrary bipartite graph, and let $d_{i}$ be the degree of a vertex $i$. Suppose that $0\leqslant x\leqslant 1$ and $y\geqslant 1$, or $0\leqslant y\leqslant 1$ and $x\geqslant 1$. Then
$\widetilde{T}_{H}(x,y)\geqslant\prod_{i\in A}\left(1+\frac{x-1}{d_{i}+1}\right)\cdot\prod_{j\in B}\left(1+\frac{y-1}{d_{j}+1}\right).$
The proof of Lemma 2.8 is based on the following inequality of Harris [12] that is also a special case of the FKG-inequality [11].
@@
Let $S_{1}\times S_{2}\times\dots\times S_{N}$ be a set, and suppose that $\mu=\mu_{1}\otimes\dots\otimes\mu_{N}$ is a product measure on it. Let $X_{1},\dots,X_{t}$ be non-negative monotone increasing functions in the sense that if $x_{i}\geqslant x_{i}^{\prime}$ for $i=1,\dots,N$, then for $1\leqslant j\leqslant t$ we have
$X_{j}(x_{1},\dots,x_{N})\geqslant X_{j}(x_{1}^{\prime},\dots,x_{N}^{\prime}).$
@@
$\mathbb{E}_{\mu}\left[\prod_{j=1}^{t}X_{j}\right]\geqslant\prod_{j=1}^{t}\mathbb{E}_{\mu}[X_{j}].$
Moreover, let $Y$ be a monotone decreasing function, that is, for $x_{i}\geqslant x_{i}^{\prime}$ $(1\leqslant i\leqslant N)$ we have
$Y(x_{1},\dots,x_{N})\leqslant Y(x_{1}^{\prime},\dots,x_{N}^{\prime}).$
If $X$ is monotone increasing and $Y$ is monotone decreasing, then
$\mathbb{E}_{\mu}[XY]\leqslant\mathbb{E}_{\mu}[X]\mathbb{E}_{\mu}[Y].$
In this paper, we will use Harris’s inequality for both the product space of $[0,1]$ intervals, that is, $0,1]^{N}$, and for the discrete set $\{0,1\}^{A}$ for some set $A$.
In what follows, we repeatedly use the following crucial idea to express $\widetilde{T}_{H}(x,y)$. We can create a random ordering of the vertices of $H$ as follows: for each vertex $i$ we choose a uniform random number $x_{i}$ from the interval $[0,1]$. The numbers $x_{i}$ then determine the ordering of the vertices of $H$. The probability that two numbers are equal is 0.
@@
$Z_{i}(x_{i},\{x_{j}\}_{j\in B})=\left\{\begin{array}[]{ll}x&\mbox{if}\ \max_{j\in N_{H}(i)}x_{j}\leqslant x_{i},\\ 1&\mbox{if}\ \max_{j\in N_{H}(i)}x_{j}>x_{i}.\end{array}\right$.